OpenStudy (hba):

If 8^y=2 and 3^x+y=81 then x= ?

OpenStudy (hba):

I got 11/3 ?

OpenStudy (hba):

Right or wrong @kropot72

OpenStudy (anonymous):

2^(3y) = 2^(1) 3y = 1 y = 1/3..

OpenStudy (hba):

I have already solved it and i got 11/3,right or wrong ?

OpenStudy (kropot72):

I get 3.996

OpenStudy (hba):

@kropot72 LoL :D

OpenStudy (hba):

You could have got 3.66667

OpenStudy (hba):

@waterineyes Yes ?

OpenStudy (hba):

So 11/3 is not the right answer ?

OpenStudy (anonymous):

3^{x+1} = 242 (x+1) log(3) = log(242) x + 1 = 4.9964 Solve this to get x..

OpenStudy (kropot72):

\[3^{x}=80\frac{2}{3}\] \[x \ln 3=\ln 80\frac{2}{3}\] \[x=\frac{\ln 80.66666}{\ln 3}=3.996\]

OpenStudy (anonymous):

It's 3.9964..

OpenStudy (anonymous):

Go with @kropot72

OpenStudy (anonymous):

If you are confident about x = 11/3 then tell us how did you get that @hba

OpenStudy (hba):

3^x+y=81 3^x+y=3^4 x+y=4 x=4-y x=4-1/3 x=12-1/3 x=11/3

OpenStudy (anonymous):

Man that x+y whole is in power???

OpenStudy (hba):

yeah ?

OpenStudy (anonymous):

Is it 3^{x+y} or 3^{x} + y..

OpenStudy (hba):

\[3^{x+y}\]

OpenStudy (anonymous):

Then you are right...

OpenStudy (hba):

Thanks Alot.

OpenStudy (anonymous):

@hba please use brackets to show properly...

OpenStudy (hba):

@waterineyes Sorry for the confusion.

OpenStudy (anonymous):

It is okay dear...

OpenStudy (kropot72):

Next time please type 3^(x+y). Thanks in advance :)